# The trichotomy of primordial black holes initial conditions

### v1

@@ Abstract
We show that the threshold to form a black hole, in an asymptotically flat and radiation dominated Friedman-Robertson-Walker (FRW) Universe, is not solely (mainly) determined by the behaviour of the compaction function at its maximum, as earlier thought, but also by the three-dimensional curvature at smaller (but super-horizon) scales, which we call “the core”. We find three classes of initial conditions characterized by an open (O), closed (C), or flat (F) FRW core surrounded by a shell with higher three-dimensional curvature. In the C case, the core helps the collapse so that the black hole formation threshold is there the lowest among all cases. Type-II black holes might only be generated by Type-O or F (each of those with different thresholds, with O being the highest) or by a Type-C with an effective F core. Finally, we argue that an F core is typically more probable for a sharp power spectrum, however, it is also more likely related to non-spherical initial conditions. On the other hand, a very broad power spectrum, which might be related to the observed NanoGrav signal, would favor the formation of Type-I black holes with a mass spectrum peaked at the Infra-Red scale.
@@ I Introduction
The knowledge of the threshold leading to black holes formation is a key ingredient in predicting their abundance in our Universe. In a seminal paper ss , Shibata and Sasaki argued that the conditions to form a black hole from a large over-density is fully determined by the behaviour of the compaction function, which is basically the gravitational potential within a sphere of radius $r$, once the infinitely long wavelength part of the perturbation is extracted. In the case of Type-I black holes, where the compaction function has only one maximum, Musco musco set the threshold to trigger gravitational collapse to the magnitude of this value. Although this threshold does depend upon the full shape of the compaction function musco , it has been argued in universal that it is mainly determined by its curvature at the maximum. With this at hands, an analytical formula, matching up to few percent error the numerical results albert , has been provided. This, opened up the doors to construct the non-linear statistics of primordial, Type-I, black holes sheth .
Type-II black holes instead are characterized by the presence of a minimum in the compaction function, surrounded by two maximums. The distinction between this and the Type-I case is mathematically determined by the magnitude at the maximum ($r_{m}$) of a variable constructed from the linear smoothed over-density multiplied by the square of the background areal radius. This variable, dubbed as $g$, gained a central role in the statistics of Primordial Black Holes (PBHs) abundances sheth . Statistically, $g$ inherit the probability distribution of curvature perturbations.
Until very recently, Type-II black holes were discarded as they needed a larger values of $g$ ($\frac{4}{3}$) than those of Type-I, and this was thought to be unlikely. Nevertheless, in a recent paper jacopo , the Authors showed that, in the case of a broad power spectrum, the non-linear statistics of Type-I profiles, developed by the use of the analytical formula of universal , prefers threshold values extremely close to Type-II amplitudes.
Shortly after the breakthrough on numerical simulation of Type-II black holes T2 , which resolved the coordinate singularity found in kopp , Escrivà in w , showed that for certain curvature profiles of the linear smoothed over-density, parameterized by their dimensionless curvature at $r_{m}$ ($w$), would produce Type-I or II black holes dependently upon the size of $w$. This shed doubts on the validity of the analytical formula universal , which, for Type-I, was supposed to be valid for any value of $w$.
@@
Consider an isolated spherically symmetric scalar perturbation. At super-horizon scales, a local patch of the perturbed Universe can be approximated by a Friedman-Robertson-Walker (FRW) metric with certain constant three-dimensional curvature (separate Universe). This patch, which we call the core, is either a closed (Type-C), an open (Type-O) or a flat (Type-F) FRW Universe. By looking at larger distances however, the spacetime becomes more and more inhomogeneous until reaching a peak of the co-moving three-dimensional curvature
@@
What we empirically found is that, while Type-O and F would not follow the analytical formula of universal for large values of $w$, as already noted in w for the F case, a Type-C fluctuation, with a non-negligible core (i.e. a core which is not effectively flat), would instead follow it increasingly better for higher values of $w$.
In other words, the three core cases provide three distinct thresholds for the maximum of $g$ being the C type threshold the lowest. The reason of it is that in the C case the core would help the collapse. Whether this extra ingredient is relevant or not, depends on the spatial extension and amplitude of the internal core, which is related a) to the sharpness of the surrounding shell parameterized by $w$ and b) by its amplitude:
@@ II Types of initial conditions
where we have defined, as in sheth , $g(r)=-\frac{4}{3}r\zeta^{\prime}(r)$. Mathematically, this variable can also be written in terms of the smoothed linear over density up to a radius $r$ ($\delta_{r}$) sheth , i.e. $g\equiv a_{i}H_{i}^{2}r^{2}\delta_{r}$, where $H_{i}$ and $a_{i}$ (set to be $1$), are the initial Hubble and scale factors. The compaction function represents the gravitational potential enclosed in a radius $r$ generated by the non-linear over-density when the background energy-density is extracted.
The maximum of $g$, which we shall name $r_{m}$, corresponds to a maximum of the compaction function for $g(r_{m})<\frac{4}{3}$ (Type-I black holes), and a minimum for $g(r_{m})>\frac{4}{3}$ (Type-II). For Type-I black holes, the threshold is typically given in terms of the maximum of the compaction function as there, the gravitational potential is the highest. In Type-II black holes, however, the minimum of the compaction function is surrounded by two maximums. The compaction function within the two maximums, being under-critical, fights against collapse, thus, the formation of a black hole is triggered whenever the magnitudes of the non-linear over-density at $r=r_{m}$ is smaller than a threshold. In terms of the linear smoothed over-density at $r_{m}$ however, both cases would require $g(r_{m})>g_{c}$, where $g_{c}$ is profile dependent musco . Nevertheless, for Type-I black holes, it has been found in universal that $g_{c}$ mostly depends on
@@ II.1 The core
According to gradient expansion, at super-horizon scales, a local patch of the spacetime is well approximated by a FRW Universe with a specific three curvature and the initial conditions for PBH formation can be classified by the sign of ${\cal R}(x\rightarrow 0)\sim 6K(x\rightarrow 0)$, as in the following table:
@@ 0 F Flat
Primordial black holes are then formed by the next to leading order inhomogeneities in the three-dimensional curvature.
@@
For Type-II profiles (6) cannot be inverted as it passes throughout an extrema. Thus, we consider instead
@@ II.2 Initial conditions
The profiles considered in w correspond to Type-F. To make our point, we have taken the exponential (Type-F) profile as a base and then deformed it in order to obtain a closed or open FRW core. Starting from the base, we then have:
@@
    valid for $\beta>2$.
@@
    is valid for $\beta>2$.
Apart from the core behaviour, which is characteristic of each profile, all three profiles present a maximum at $r=r_{p}$ and a subsequent minimum. The radius $r_{m}$, discussed earlier, sits in between (see e.g. Fig.1). The similarities of the curves are due to the fact that we are considering only a single peak $\zeta$ profiles, the most luckily configuration for PBHs.
@@ III Numerics
To track the black hole formation more efficiently, we have modified the code including an auxiliary function defined as twice the Misner-Sharp mass $M$ ms over the areal radius $R$
@@
An apparent horizon leading to a black hole is only formed when $\mathcal{C}_{0}$, at its maximum, reaches $1$ twice. The first $1$ would signal the cosmological horizon crossing of the super-horizon perturbation. After this, because of the dilution of the Universe expansion, $\mathcal{C}_{0}$ would decrease. However, over threshold, the decay would then be followed by a growth and a sub-horizon trapped surface would form whenever $\mathcal{C}_{0}\sim 1$ for the second time.
Large values of $w$ (>30>30), i.e. initial sharp profiles, required a substantial modification of the number of Chebyshev points ($N_{\text{cheb}}$) in the pseudo-spectral method used in w . The code becomes increasingly unstable for $w>50$ in the Type-C case. However, as we shall show, the threshold already saturates at those values with no need of extra probes [^2] In fact, Type-C configurations for $w>50$ when simulated with the SPriBHoS-II code, show a clear trend to the formation of a trapped horizon with a saturated threshold. This happens way before the code breaks because of Gibbs phenomena..
@@ IV Thresholds
Here we present the thresholds values, $g_{c}$, for different profiles. In Fig.2, we can clearly see that as $w$ increases, the trichotomy appears. For Type-F and Type-C perturbations the crossing between Type-I and Type-II happens around $w\sim 34$, while for Type-O happens before, at $w\sim 29$. We also see the trend anticipated earlier that the lowest thresholds are related to Type-C where the cores help the collapse. On the other hand, the highest thresholds are related to Type-O where instead the cores fight the collapse.
For large values of $w$, Type-O and Type-F become Type-II, as already noted in w . On the other hand, Type-C perturbations lie in both Type-I and Type-II regions. We will explore this peculiar behaviour in the following sections.
Figure 2: Numerical threshold, in terms of $g$, for the three types of perturbations. The dashed line correspond to $g=\frac{4}{3}.$ All the points correspond to different configurations of the $\alpha$ and $\beta$ parameters (see Appendix A).
@@
As we already mentioned, we shall work instead with $g$ which is more convenient for Type-II perturbations. Notice however that, because the relation between $\mathcal{C}$ and $g$ is quadratic, the error in terms of $g$ is generically higher than that of $\mathcal{C}$ (see Fig.4).
In Fig. 3 we selected, from Fig.2, all Type-I configurations.
@@
There, the relative deviation between the analytical threshold and the numerical one, $d\equiv|\delta_{N}-\delta_{A}|/\delta_{N}$, is $\sim 5\%$ (which reduces to $2\%$ for $\mathcal{C}_{c}$) for values of $w\sim 16$ and decreases to $\sim 1.8\%$, as $w$ increases ($0.15\%$ for $\mathcal{C}_{c}$), c.f. Fig.4.
Type-II perturbations however require a higher threshold as there the core would not help for the collapse of the external (higher density) shell. Illustrative examples are given in Fig.5, where Type-II is obtained by lowering the core from a Type-I(C).
Figure 5: ${\cal R}(r)$ for the three types of perturbation at threshold ($w\simeq 38.6$).
In w it was shown that, at order $w\sim 30$ all the initial profiles considered (that can be check to be of Type-F), at threshold, would produce Type-II black holes. In Fig.3, we have explicitly shown that this is not always the case for Type-C, while it is true for Type-F and O. Nevertheless, there is also a distinction between those two. Considering the fitting formula proposed in w , i.e.
@@
for $w\gtrsim 30$, one can see that it would only work for Type-F. Testing (18) for flat profiles (8) lead to a relative deviation with respect to numeric of less than $0.3\%$. However, for Type-O, this analytical formula does not fit so well the numerical values, leading to a relative deviation of $\sim 5\%$. For those Type-C profiles that cross to Type-II region, the proposed analytical formula gives us a full range of values for similar values of $w$, from $d\sim 10\%$, which clearly largely deviate from the analytical formula, to $d\sim 2\%$ (see Fig.7).
Figure 6: Numerical threshold, in terms of $g$, for Type-II perturbations.
Figure 7: Relative deviation, $d$, in terms of $g$ for Type-II perturbations using (18).
@@ V Transition between closed and flat perturbations
As we have shown, Type-C perturbations are characterized by constant strictly positive core values for ${\cal R}(x)$ or ${\cal R}(r)$ (Fig.1). One may ask what happens if this constant value decreases approaching zero, if it would still behave as Type-C or rather as a Type-F [^3] We thank Jaume Garriga for posing this question..
As we have argued, if the core corresponds to a closed FRW, it helps the collapse of the perturbation. In this case, a “low” threshold is in order, leading to a Type-I BH. On the other hand, in the flat case, the core does not help the collapse, implying a higher threshold leading to Type-II BHs.
[…]
@@
In Fig.8, we can see that, fixing $w_{\cal R}$ and starting from a Type-C(I) perturbation, we can go to a Type-II by lowering the core amplitude. Once the perturbation becomes Type-II, we say that the core behave as an effective flat one.
@@
  • The amplitude of the core is large enough with respect to the peak amplitude.
  • The peak is sharp enough.
The first condition is actually connected to the second one and can be tested computing the ratio
@@
Figure 9: Values for ${\cal Q}\equiv\frac{{\cal R}(0)}{{\cal R}(r_{p})}\Big|_{c}$ for different values of $w_{\cal R}$. The blue points correspond to Type-I perturbations, while the red ones to Type-II. The line correspond to a tentative transition region (Eq.(21)) that depends on the value of $w_{\cal R}$.
If the peak is broad enough (i.e. $w_{\cal R}$ small), we can consider it as a closed FRW, which will help the collapse. That’s the reason why for small values of $w_{\cal R}$ the three types of perturbations are Type-I. On the other hand, for higher values of $w_{\cal R}$ the peak gets thinner and loses relevance with respect to the core.
@@
Once we have identified the region of effective Type-F, we can interpret the Type-C(II) points in Fig.2 as those below the curve $\cal Q$ (c.f. Fig.10).
Figure 10: Numerical threshold, in terms of $g$ as a function of $w$, for the three types of perturbations, including the effective flat profiles. The dashed line correspond to $g=\frac{4}{3}.$
@@
Figure 11: Numerical threshold, in terms of $g$ as a function of $w_{\cal R}$, for the three types of perturbations, including the effective flat profiles. The dashed line correspond to $g=\frac{4}{3}.$

### v2

@@ Abstract
We show that the threshold to form a black hole, in an asymptotically flat and radiation-dominated Friedman-Robertson-Walker (FRW) Universe, is not solely (mainly) determined by the behaviour of the compaction function at its extrema, as earlier thought, but also by the Ricci scalar of the spatial geometry at smaller (but super-horizon) scales, which we call “the core”. We introduce three classes of initial conditions characterised by an open (O), closed (C), or flat (F) FRW core surrounded by a shell with higher three-dimensional curvature. In the C case, the core helps the collapse so that the black hole formation threshold is the lowest among all cases. Type-II black holes might only be generated by Type-O or F (each of those with different thresholds, with O being the highest) or by a Type-C with a negligible Ricci scalar at the centre, which we call an effective F core. Finally, we argue that an F core is typically more probable for a sharp power spectrum, however, it is also more likely related to non-spherical initial conditions. On the other hand, a very broad power spectrum, which might be related to the observed NanoGrav signal, would favour the formation of Type-I black holes with a mass spectrum peaked at the Infra-Red scale.
@@ I Introduction
The knowledge of the threshold leading to black hole formation is a key ingredient in predicting their abundance in our Universe. In a seminal paper ss , Shibata and Sasaki argued that the conditions to form a black hole from a large over-density is fully determined by the behaviour of the compaction function, which is basically the gravitational potential within a sphere of radius $r$, once the infinitely long wavelength part of the perturbation is extracted. In the case of Type-I black holes, where the compaction function has only one maximum, Musco musco set the threshold to trigger gravitational collapse to the magnitude of this value. Although this threshold does depend upon the full shape of the compaction function musco , it has been argued in universal that it is mainly determined by its curvature at the maximum. With this at hand, an analytical formula, matching up to a few percent error, the numerical results albert , has been provided. This opened up the doors to construct the non-linear statistics of primordial, Type-I, black holes sheth .
Type-II black holes instead, are characterised by the presence of a minimum in the compaction function, surrounded by two maxima. The distinction between this and the Type-I case is mathematically determined by the magnitude at the maximum ($r_{m}$) of a variable constructed from the linear smoothed over-density multiplied by the square of the background areal radius. This variable, dubbed as $g$, gained a central role in the statistics of Primordial Black Holes (PBHs) abundances sheth . Statistically, $g$ inherits the probability distribution of curvature perturbations.
Until very recently, Type-II black holes were discarded as they needed larger values of $g$ ($\frac{4}{3}$) than those of Type-I, and this was thought to be unlikely. Nevertheless, in a recent paper jacopo , the Authors showed that, in the case of a broad power spectrum, the non-linear statistics of Type-I profiles, developed by the use of the analytical formula of universal , prefer threshold values extremely close to Type-II amplitudes.
Shortly after the breakthrough on numerical simulation of Type-II black holes T2 , which resolved the coordinate singularity found in kopp , Escrivà in w , showed that for certain curvature profiles of the linear smoothed over-density, parametrised by their dimensionless curvature at $r_{m}$ ($w$), would produce Type-I or II black holes dependently upon the size of $w$. This shed doubts on the validity of the analytical formula universal , which, for Type-I, was supposed to be valid for any value of $w$.
@@
Consider an isolated spherically symmetric scalar perturbation. At super-horizon scales, a local patch of the perturbed Universe can be approximated by a Friedman-Robertson-Walker (FRW) metric with certain constant three-dimensional curvature (separate Universe). This patch, which we call the core, is either a closed (Type-C), an open (Type-O) or a flat (Type-F) FRW Universe. By looking at larger distances, however, the spacetime becomes more and more inhomogeneous until reaching a peak of the co-moving three-dimensional curvature
@@
What we empirically found is that, while Type-O and F would not follow the analytical formula of universal for large values of $w$, as already noted in w for the F case, while a Type-C fluctuation, with a non-negligible core, would instead follow it increasingly better for higher values of $w$.
In other words, the three core cases provide three distinct thresholds for the maximum of $g$, with the C type threshold being the lowest. The reason of it is that in the C case, the core geometry would help the collapse. Whether this extra ingredient is relevant or not depends on the spatial extension and amplitude of the internal core, which is related a) to the sharpness of the surrounding shell parametrised by $w$ and b) by its amplitude:
@@ II Types of initial conditions
where we have defined, as in sheth , $g(r)=-\frac{4}{3}r\zeta^{\prime}(r)$, $r=0$ is the locus of the over-density peak (for a maximum) or trough (for a minimum). Mathematically, this variable can also be written in terms of the smoothed linear over density up to a radius $r$ ($\delta_{r}$) sheth , i.e. $g\equiv a_{i}H_{i}^{2}r^{2}\delta_{r}$, where $H_{i}$ and $a_{i}$ (set to be $1$), are the initial Hubble and scale factors. The compaction function represents the gravitational potential enclosed in a radius $r$ generated by the non-linear over-density when the background energy-density is extracted.
The maximum of $g$, which we shall name $r_{m}$, corresponds to a maximum of the compaction function for $g(r_{m})<\frac{4}{3}$ (Type-I black holes), and a minimum for $g(r_{m})>\frac{4}{3}$ (Type-II). For Type-I black holes, the threshold is typically given in terms of the maximum of the compaction function, as there, the gravitational potential is the highest. In Type-II black holes, however, the minimum of the compaction function is surrounded by two maxima. The compaction function within the two maximums, being under-critical, fights against collapse, thus, the formation of a black hole is triggered whenever the magnitude of the non-linear over-density at $r=r_{m}$ is smaller than a threshold. In terms of the linear smoothed over-density at $r_{m}$, however, both cases would require $g(r_{m})>g_{c}$, where $g_{c}$ is profile dependent musco .
Nevertheless, for Type-I black holes, it has been found in universal that $g_{c}$ mostly depends on the physics around the collapsing shell surrounding the largest over-density value (the shell from now on). Specifically, the threshold has been found to depend upon the curvature of the smoothed linear over-density at its maximum, i.e., it would only depend on
@@
On the contrary, and this is what we are going to argue, in Type-II configurations, the central value of the over-density plays a non-trivial role for the threshold.
@@ II.1 The core
According to gradient expansion, at super-horizon scales, a local patch of the spacetime is well approximated by a FRW Universe with a specific three-dimensional curvature, and the initial conditions for PBH formation can be classified by the sign of ${\cal R}(x\rightarrow 0)\sim 6K(x\rightarrow 0)$, as in the following table:
@@ 0 F Flat
Primordial black holes are then formed by the next-to-leading order inhomogeneities in the three-dimensional curvature.
@@
For Type-II profiles, (6) cannot be inverted as it passes through an extrema. Thus, we consider instead
@@
As we can already anticipate, the amplitude and size of the core can either help or resist the collapse of the shell. It indeed turns out, as we are going to see, that Type-II configurations are related to the case in which the core cannot be forgotten in the computation of the threshold.
@@ II.2 Initial conditions
The profiles considered in w , which for large values of $w$ implied Type-II configurations, correspond to Type-F. To make our point, we have taken the exponential (Type-F) profile as a base and then deformed it in order to obtain a closed or open FRW core. In this sense, what we are going to consider are collapsing shells characterised by the three possible configurations for the cores and connect those with Type-I and II black holes.
Starting from the base, we then consider:
@@
    valid for $\beta\geq 2$. This is a physical condition, as for smaller $\beta$ the Poisson equation would diverge at the centre.
@@
    is valid for $\beta\geq 2$.
[…]
@@ III Numerics
To track the black hole formation more efficiently, we have modified the code, including an auxiliary function defined as twice the Misner-Sharp mass $M$ ms over the areal radius $R$
@@
An apparent horizon leading to a black hole is only formed when $\mathcal{C}_{0}$, at its maximum, reaches $1$ twice. The first $1$ would signal the cosmological horizon crossing of the super-horizon perturbation. After this, because of the dilution of the Universe expansion, $\mathcal{C}_{0}$ would decrease. However, over threshold, the decay would then be followed by a growth, and a sub-horizon trapped surface would form whenever $\mathcal{C}_{0}\sim 1$ for the second time.
Large values of $w$ (>30>30), i.e. initial sharp profiles, required a substantial modification of the number of Chebyshev points ($N_{\text{cheb}}$) in the pseudo-spectral method used in w . The code becomes increasingly unstable for $w>50$ in the Type-C case. However, as we shall show, the threshold already saturates at those values with no need of extra probes [^2] In fact, Type-C configurations for $w>50$ when simulated with the SPriBHoS-II code, show a clear trend to the formation of a trapped horizon with a saturated threshold. This happens way before the simulation breaks because of Gibbs phenomena..
@@ IV Thresholds
Here we present the threshold values, $g_{c}$, for different profiles. In Fig.2, we can clearly see that as $w$ increases, the trichotomy appears. For Type-F and Type-C perturbations, the crossing between Type-I and Type-II happens around $w\sim 34$, while for Type-O happens before, at $w\sim 29$. We also see the trend anticipated earlier, that the lowest thresholds are related to Type-C, where the cores help the collapse. On the other hand, the highest thresholds are related to Type-O, where instead the cores fight the collapse.
This is one of the main results of our paper, namely that for large enough cores, the threshold cannot only depend upon the shell structure.
Figure 2: Numerical threshold, in terms of $g$, for the three types of perturbations. The dashed line corresponds to $g=\frac{4}{3}.$ All the points correspond to different configurations of the $\alpha$ and $\beta$ parameters (see Appendix A).
For large values of $w$ (still in Fig. 2), Type-O and Type-F become Type-II, as already noted in w . On the other hand, Type-C perturbations lie in both Type-I and Type-II regions. We will explore this peculiar behaviour in the following sections, nevertheless, we can already anticipate that Type-C should smoothly behave as Type-F for small core amplitudes. This is precisely what happens in the Fig. 2.
IV.1 Type-I
@@
As we already mentioned, we shall work instead with $g$, which is more convenient for Type-II perturbations. Notice, however, that because the relation between $\mathcal{C}$ and $g$ is quadratic (4), the error in terms of $g$ is generically higher than that of $\mathcal{C}$ (see Fig.4).
In Fig. 3, we selected, from Fig.2, all Type-I configurations.
@@
There, the relative deviation between the analytical threshold ($\delta_{A}$) and the numerical one ($\delta_{N}$), $d\equiv|\delta_{N}-\delta_{A}|/\delta_{N}$, is $\sim 5\%$ (which reduces to $2\%$ for $\mathcal{C}_{c}$) for values of $w\sim 16$ and decreases to $\sim 1.8\%$, as $w$ increases ($0.15\%$ for $\mathcal{C}_{c}$), c.f. Fig.4.
IV.2 Type-II
Type-II perturbations require a higher threshold with respect to those of Type-I, as there the core would not help for the collapse of the external (higher density) shell. Illustrative examples are given in Fig.1, where Type-II is obtained by lowering the core from a Type-C(I).
In w it was shown that, at order $w\sim 30$ all the initial profiles considered (that can be checked to be of Type-F), at threshold, would produce Type-II black holes. In Fig.3, we have explicitly shown that this is not always the case for Type-C, while it is true for Type-F and O. Nevertheless, there is also a distinction between those two. Considering the fitting formula proposed in w , i.e.
@@ IV.2 Type-II
for $w\gtrsim 30$, one can see that it would only work for Type-F. Testing (18) for flat profiles (8) leads to a relative deviation with respect to the numeric of less than $0.3\%$. However, for Type-O, this analytical formula does not fit so well the numerical values, leading to a relative deviation of $\sim 5\%$. For those Type-C profiles that cross to the Type-II region, the proposed analytical formula gives us a full range of values for similar values of $w$, from $d\sim 10\%$, which clearly largely deviate from the analytical formula, to $d\sim 2\%$ (see Fig.5).
Figure 5: The left panel is the numerical threshold, in terms of $g$, for Type-II perturbations. The right panel is the relative deviation, $d$, in terms of $g$ for Type-II perturbations using (18).
@@ V Transition between closed and flat perturbations
As we have shown, Type-C perturbations are characterised by constant strictly positive core values for ${\cal R}(x)$ or ${\cal R}(r)$ (Fig.1).
One may ask what happens if this constant value decreases approaching zero, if it would still behave as Type-C or rather as a Type-F [^3] We thank Jaume Garriga for posing this question..
As we have argued, the trichotomy appears as the shell gets thinner, and the core plays a role. Then, if the core corresponds to a closed FRW, it helps the collapse of the perturbation. In this case, a “low” threshold is in order, leading to a Type-I BH. On the other hand, in the flat case, the core does not help the collapse, implying a higher threshold leading to Type-II BHs.
Because of the central importance of the size of the collapsing shell, we propose a new dimensionless parameter:
@@
In Fig.6, we can see that, fixing $w_{\cal R}$ we can go to a Type-II by lowering the core amplitude from a Type-I configuration with closed core (which we call Type-C(I)). Once the perturbation becomes Type-II, which requires a higher threshold, the core effect is negligible, and hence, it effectively acts as a flat one, we call this Type-C(II) configuration.
@@
  • The amplitude of the core, ${\cal R}(0)$, is large enough with respect to the peak amplitude, ${\cal R}(r_{p})$,.
  • The peak is sharp enough, meaning $w_{\cal R}$ is large enough.
The first condition is actually connected to the second one and can be tested by computing the ratio
@@
Figure 7: Values for ${\cal Q}\equiv\frac{{\cal R}(0)}{{\cal R}(r_{p})}\Big|_{c}$ for different values of $w_{\cal R}$. The blue points correspond to Type-I perturbations, while the red ones to Type-II. The line corresponds to a tentative transition region (Eq.(21)) that depends on the value of $w_{\cal R}$.All the data points correspond to the different profiles shown in Appendix A.
If the peak is broad enough (i.e. $w_{\cal R}$ small), we can consider it as a closed FRW, which will help the collapse. Then, the core behaviour is not noticeable. That is the reason why, for small values of $w_{\cal R}$ the three types of perturbations are Type-I. On the other hand, for higher values of $w_{\cal R}$, the peak gets thinner and loses relevance with respect to the core.
@@
Once we have identified the region of Type-C perturbations with a negligible core, we can interpret the Type-C(II) points as effective Type-F ones (those below the curve $\cal Q$, c.f. Fig.8).
Figure 8: Numerical threshold, in terms of $g$ as a function of $w$, for the three types of perturbations, including the Type-C(II) profiles that behave effectively as flat. The dashed line corresponds to $g=\frac{4}{3}$. All the data points correspond to the different profiles shown in Appendix A.
@@
Figure 9: Numerical threshold, in terms of $g$ as a function of $w_{\cal R}$, for the three types of perturbations, including the effective flat profiles. The dashed line corresponds to $g=\frac{4}{3}.$